two angles are complementary the larger angle is 5 less than 4 times than the smaller angle
Answer by nerdybill(7384)

two angles are complementary the larger angle is 5 less than 4 times than the smaller angle
.
Let x = smaller angle
then
4x-5 = larger angle
.
x + 4x-5 = 90
5x-5 = 90
5x = 95
x = 19 degrees (smaller angle)
.
larger angle:
4x-5 = 4(19)-5 = 76-5 = 71 degrees
.
answer: 71 and 19 degrees